Two railway stations are at a distance of 96 km from each other. One train covers this distance in 40 minutes less than the other. The speed of the first train is 12 km/h faster than that of the second. Determine the speeds of both the trains.
Answer by Alan3354(69443)

Two railway stations are at a distance of 96 km from each other. One train covers this distance in 40 minutes less than the other. The speed of the first train is 12 km/h faster than that of the second. Determine the speeds of both the trains.
---------------
For the slower train, t = d/r = 96/r
For the faster, t- 2/3 = 96/(r+12)
---
96/r - 2/3 = 96/(r+12)
Multiply by 3r
288 - 2r = 96*3r/(r+12)
144 - r = 144r/(r+12)
(144-r)*(r+12) = 144r
-r^2 + 132r + 1728 = 144r
-r^2 - 12r + 1728 = 0
r^2 + 12r - 1728 = 0
(r+48)*(r-36) = 0
---------------
Slow train goes 36 km/hr
Fast train goes 48 km/hr